OpenStudy (anonymous):

solve x^4-109x^2+900=0 a)3,-3,10,-10 this is the correct answer I just don't know how to do the problem work shown would be a huge help

OpenStudy (nikato):

Do u know the box and cross method

OpenStudy (anonymous):

yep

OpenStudy (nikato):

So in ur problem, u are trying to look for two factors of 900 that also add up to -109

OpenStudy (anonymous):

ok

OpenStudy (nikato):

Do u know what they are?

OpenStudy (anonymous):

umm no

OpenStudy (nikato):

They are -9 and -100 -9(-100)=-900 -9+-100=-109

OpenStudy (nikato):

So the factored form would be (x^2-9)(x^2-100)

OpenStudy (nikato):

And u can simplify that even more (X-3)(x+3)(x-10)(x+10)

OpenStudy (nikato):

And u should b able to find x. Are u ok with that^

OpenStudy (anonymous):

yess thank you sooo much that is a huge help